A ball is thrown horizontally from the top of a 80 m high cliff. If the initial horizontal velocity is 10.0 m/s, find
(ii) how far from the foot of the cliff it strikes the ground
1
Expert's answer
2020-07-02T17:14:47-0400
"x=vt=v\sqrt{\frac{2h}{g}}\\x=10\sqrt{\frac{2(80)}{9.8}}=40.4\ m"